# Triplet of kindred prompt-2 emitters in mass-8 proton-rich nuclei

### v1

@@ Abstract
In this work we consider the nature of the second peak in Fig. 2 at =8.4 MeV. This peak is associated with a state of lower excitation energy than that for the narrow 0 peak as it decays to the ground state of 6Li [CharityLong:2025], i.e. there is no coincidence ray and thus =. The existence of this state was confirmed from a second experiment, where the fragmentation products from an =65-MeV 13O beam with a 9Be target [CharityLong:2025] are analyzed with the invariant-mass method. The weighted mean resonance parameters, from the two data sets, are =8.40(4) MeV and =0.88(7) MeV [CharityLong:2025].
@@
Figures 4(b) and 4(c) show the decay correlations for the three states of interest in terms of the fraction of the decay energy in the - and -core sub-systems. (The core here refers to the daughter, i.e., 6Li[1+, T = 0], 6Li[0+, T = 1], and 6Be[0+, T = 1], from bottom to top in Fig. 3.) The results for 8C(0) and 8B(0) (red and green data points) are from [Charity:2011, Brown:2014]. Background-subtracted correlations for the 8.4-MeV state were extracted in Ref. [CharityLong:2025] from the two invariant-mass data sets (with the 9C and 13O beams) and found to be consistent. The results shown by the blue data points represents the average of the two data sets with the error bars indicating their spread.
@@
For the data set with the 9C beam, an invariant-mass peak for +3He+ exit channel was found with =8.23(18) MeV and =1.30(33) [CharityLong:2025] which are close to the values for the 8.4-MeV peak. This could possibly represents a second decay branch of this 2 emitter. If so, the +3He+ branch has a yield 1.2(4) times that of the 2 branch [CharityLong:2025]. Correlations indicate that the +3He+ exit channel proceeds by an initial 3He+5Li decay and possibly a smaller contribution from +7Be(5/2) decay [CharityLong:2025].
@@
Discussion. — Based on the SA-NCSM calculations, candidate levels that can be produced via -wave proton knockout from the 9C beam (0+, 1+, 2+, and 3+) and located in excitation close to the newly observed 2 emitter are the 1, 2 and the 3 states [CharityLong:2025]. These calculations largely reproduce the known levels at lower excitation. The single-proton decay energies of these states to the ground (3/2-) and first excited (1/2-) states in 7Be are more than twice the height of the Coulomb plus centrifugal barrier. In order for a 2 decay branch to complete, the 1 channels must be associated with very small spectroscopic factors. The SA-NCSM spectroscopic factors for the candidate states to these low-lying states in 7Be (Table 1) are indeed quite small (.
The model calculations indicate that the three candidates have significant spectroscopic factors for proton decay to either of the two higher-lying adjacent 5/2- states in 7Be [CharityLong:2025]. Experimentally, the 5/2 state decays almost exclusively to the 3He+ channel [Tombrello:1963, Spiger:1967] while the 5/2 state decays predominantly to the +6Li(g.s.) channel [McCray:1963] which is consistent with the SA-NCSM calculations [CharityLong:2025] and recent Gamow shell-model results [Fernandez:2023].
@@
The possibility that the 1 state has a significant decay branch to the +3He+ exit channel is not supported by the SA-NCSM calculations for which the 3He+5Li and +7Be(5/2) decay branches are minor [CharityLong:2025].
The same structure should be present in the levels of the mirror nuclei. The mirror 1 level in 8Li (Fig. 1) has been studied via -delay triton emission of 8He. Borge et al. [Borge:1993] measured the energy spectrum, which was fit with a +5He++ sequential decay model, giving an excitation energy of 9.3(1) MeV. This decay path is the mirror of the 3He+5Li path which dominates the state observed in the +3He+ channel observed at almost the same energy. This adds support to the contention that we observed the 3He+5Li decay branch of 8B(1), however, there are other interpretations of the -energy spectrum. Barker’s analysis, allowing for the possibility of more levels and channels, deduced E∗=9.02-9.92 MeV [Barker:1996]. Grigorenko et al. using a three-body ++ model showed that the triton spectrum could be fit without introducing correlations, i.e. with a simple phase-space approach, and found E∗=8.5-8.8 MeV [Grigorenko:1996]. Clearly the deduced excitation energy, of the 1 level in the mirror 8Li, is model dependent. The SA-NCSM and the Greens function Monte Carlo calculations [Wiringa:2000] favor excitation energies at the lower end of the range of these three analyses [CharityLong:2025].
@@
Conclusions. — The nature of the =8.4-MeV resonance state in 8B observed in [CharityLong:2025] that decays to the 2+6Li channel has been established and shown to have an unexpected resemblance to neighboring resonances. Momentum correlations between the decay fragments indicate that this is a prompt 2 emitting state. Based on comparison to ab initio symmetry-adapted no-core shell-model calculations, this level is assigned as the forth 1+ state which has a spatial structure very similar to 8C and its isobaric analog state in 8B (=2 analog states). Indeed, this newly found state and the two just mentioned constitute a triplet of kindred 2 emitters with similar correlations between the decay products and which, via these decays, remove one unit of isospin, leave the spin and parity unchanged, and converts an oblate parent into a prolate daughter.

### v2

@@ Abstract
In this work we consider the nature of the second peak in Fig. 2 at =8.4 MeV. This peak is associated with a state of lower excitation energy than that for the narrow 0 peak as it decays to the ground state of 6Li [CharityLongArxiv], i.e. there is no coincidence ray and thus =. The existence of this state was confirmed from a second experiment, where the fragmentation products from an =65-MeV 13O beam with a 9Be target [CharityLongArxiv] are analyzed with the invariant-mass method. The weighted mean resonance parameters, from the two data sets, are =8.40(4) MeV and =0.88(7) MeV [CharityLongArxiv].
@@
Figures 4(b) and 4(c) show the decay correlations for the three states of interest in terms of the fraction of the decay energy in the - and -core sub-systems. (The core here refers to the daughter, i.e., 6Li[1+, T = 0], 6Li[0+, T = 1], and 6Be[0+, T = 1], from bottom to top in Fig. 3.) The results for 8C(0) and 8B(0) (red and green data points) are from [Charity:2011, Brown:2014]. Background-subtracted correlations for the 8.4-MeV state were extracted in Ref. [CharityLongArxiv] from the two invariant-mass data sets (with the 9C and 13O beams) and found to be consistent. The results shown by the blue data points represents the average of the two data sets with the error bars indicating their spread.
@@
For the data set with the 9C beam, an invariant-mass peak for +3He+ exit channel was found with =8.23(18) MeV and =1.30(33) [CharityLongArxiv] which are close to the values for the 8.4-MeV peak. This could possibly represents a second decay branch of this 2 emitter. If so, the +3He+ branch has a yield 1.2(4) times that of the 2 branch [CharityLongArxiv]. Correlations indicate that the +3He+ exit channel proceeds by an initial 3He+5Li decay and possibly a smaller contribution from +7Be(5/2) decay [CharityLongArxiv].
@@
Discussion. — Based on the SA-NCSM calculations, candidate levels that can be produced via -wave proton knockout from the 9C beam (0+, 1+, 2+, and 3+) and located in excitation close to the newly observed 2 emitter are the 1, 2 and the 3 states [CharityLongArxiv]. These calculations largely reproduce the known levels at lower excitation. The single-proton decay energies of these states to the ground (3/2-) and first excited (1/2-) states in 7Be are more than twice the height of the Coulomb plus centrifugal barrier. In order for a 2 decay branch to complete, the 1 channels must be associated with very small spectroscopic factors. The SA-NCSM spectroscopic factors for the candidate states to these low-lying states in 7Be (Table 1) are indeed quite small (.
The model calculations indicate that the three candidates have significant spectroscopic factors for proton decay to either of the two higher-lying adjacent 5/2- states in 7Be [CharityLongArxiv]. Experimentally, the 5/2 state decays almost exclusively to the 3He+ channel [Tombrello:1963, Spiger:1967] while the 5/2 state decays predominantly to the +6Li(g.s.) channel [McCray:1963] which is consistent with the SA-NCSM calculations [CharityLongArxiv] and recent Gamow shell-model results [Fernandez:2023].
@@
The possibility that the 1 state has a significant decay branch to the +3He+ exit channel is not supported by the SA-NCSM calculations for which the 3He+5Li and +7Be(5/2) decay branches are minor [CharityLongArxiv].
The same structure should be present in the levels of the mirror nuclei. The mirror 1 level in 8Li (Fig. 1) has been studied via -delay triton emission of 8He. Borge et al. [Borge:1993] measured the energy spectrum, which was fit with a +5He++ sequential decay model, giving an excitation energy of 9.3(1) MeV. This decay path is the mirror of the 3He+5Li path which dominates the state observed in the +3He+ channel observed at almost the same energy. This adds support to the contention that we observed the 3He+5Li decay branch of 8B(1), however, there are other interpretations of the -energy spectrum. Barker’s analysis, allowing for the possibility of more levels and channels, deduced E∗=9.02-9.92 MeV [Barker:1996]. Grigorenko et al. using a three-body ++ model showed that the triton spectrum could be fit without introducing correlations, i.e. with a simple phase-space approach, and found E∗=8.5-8.8 MeV [Grigorenko:1996]. Clearly the deduced excitation energy, of the 1 level in the mirror 8Li, is model dependent. The SA-NCSM and the Greens function Monte Carlo calculations [Wiringa:2000] favor excitation energies at the lower end of the range of these three analyses [CharityLongArxiv].
@@
Conclusions. — The nature of the =8.4-MeV resonance state in 8B observed in [CharityLongArxiv] that decays to the 2+6Li channel has been established and shown to have an unexpected resemblance to neighboring resonances. Momentum correlations between the decay fragments indicate that this is a prompt 2 emitting state. Based on comparison to ab initio symmetry-adapted no-core shell-model calculations, this level is assigned as the forth 1+ state which has a spatial structure very similar to 8C and its isobaric analog state in 8B (=2 analog states). Indeed, this newly found state and the two just mentioned constitute a triplet of kindred 2 emitters with similar correlations between the decay products and which, via these decays, remove one unit of isospin, leave the spin and parity unchanged, and converts an oblate parent into a prolate daughter.